# EarthDial: Turning Multi-sensory Earth Observations to Interactive Dialogues

### v1

@@ Abstract
Automated analysis of vast Earth observation data via interactive Vision-Language Models (VLMs) can unlock new opportunities for environmental monitoring, disaster response, and resource management. Existing generic VLMs do not perform well on Remote Sensing data, while the recent Geo-spatial VLMs remain restricted to a fixed resolution and few sensor modalities. In this paper, we introduce EarthDial, a conversational assistant specifically designed for Earth Observation (EO) data, transforming complex, multi-sensory Earth observations into interactive, natural language dialogues. EarthDial supports multi-spectral, multi-temporal, and multi-resolution imagery, enabling a wide range of remote sensing tasks, including classification, detection, captioning, question answering, visual reasoning, and visual grounding. To achieve this, we introduce an extensive instruction tuning dataset comprising over 11.11M instruction pairs covering RGB, Synthetic Aperture Radar (SAR), and multispectral modalities such as Near-Infrared (NIR) and infrared. Furthermore, EarthDial handles bi-temporal and multi-temporal sequence analysis for applications like change detection. Our extensive experimental results on 43 downstream applications demonstrate that EarthDial outperforms existing generic and domain-specific models, achieving better generalization across various EO tasks.
@@ 1 Introduction
  • Experimental results demonstrate that EarthDial performs well in comparison to existing domain-specific VLMs, achieving higher accuracies and better generalization across 37 downstream EO tasks.
@@ 2 Related Work
RS Instruction Datasets: A large number of instruction-following datasets have been introduced to train the RS-VLMs effectively. For example, GeoChat-Instruct [28], SkyEye-968k [67], and RS5M [71] provide extensive RS image-text pairs, supporting instruction-based VLM training on optical data. These existing datasets limit the model’s capabilities across different sensor modalities. On the other hand, MLLMs [70], provide image-text pairs from optical, SAR, and infrared images. Regardless of the large scale, the aforementioned datasets often lack diverse RS applications across various modalities, including multi-resolution, multi-spectral, and multi-temporal RS sensor data. The recently introduced dataset aims to enhance content richness with factual and deceptive questions, enabling VLMs to address a wider range of tasks in the RS domain [46]. However, there is no large-scale unified instruction-following dataset that can encapsulate the distinctive contextual complexities of diverse RS applications across different modalities, multi-resolution, multi-spectral, and multi-temporal RS sensor data. Our work is an effort to bridge this gap with an 11M instruction set to seamlessly integrate multiple earth observation modalities covering diverse spectra and with time-series imaging data for diverse RS applications.
@@ 3.2 Stage-wise Model Training
We utilized the pretrained weights from the previous stage 2 and fine-tuned the MLP and LLM layers, while keeping the ViT parameters frozen from Stage 1. The same ViT+MLP+LLM architecture was employed, with the addition of the data fusion module (Sec. 3.1) to integrate information from multiple spectral channels. This training stage allows the model to handle a broader range of tasks, including land cover classification, species detection, Methane plume detection, UHI and SAR-based ship detection.
@@ 4 EarthDial-Instruct Dataset
Pretraining Instruction Data: With EarthDial, our primary aim is to improve generalization performance on diverse downstream tasks, covering a wide range of modalities, multi-resolution, and multi-temporal data. Therefore, we curate high-quality pre-train question-answer (QA) instruction pairs from SkyScript [58] and SatlasPretrain [6] data, which includes Sentinel-2 (S2), Sentinel-1 (SAR), NAIP, and Landsat imagery along with labels. We choose InternLM-XComposer2 [22] (leading open VLM at the time of selection) for generating instructions using labels. Our data curation process involves filtering to ensure data quality. First, we filter out samples with sparse labels ($<$3). Second, we apply luminance and coverage-based filtering to remove cloudy and low spatial coverage images. Third, we prompt the InternLM-XComposer2 to generate QA instruction pairs based on the key attributes (points, polygons, object category, and position) specified in the inputs and labels. The details can be found in the supplementary material. The curated instruction stats across different imagery sources are presented in Tab. 2.
@@
Downstream Tasks Image-text Instruction: While pre-training focuses on enhancing generalization capabilities, we also need task-specific fine-tuning with diverse data types to improve downstream performance as shown in Fig. 4. To handle this, we carefully curate a large number of instruction-following datasets that cover ten diverse downstream tasks (e.g., scene classification, object detection, visual question answering, image captioning, change detection, Methane plume detection, tree species classification, local climate zones, urban heat islands, and disaster assessment), six visual modalities (Optical, SAR, S2, Infrared, NIR, and Hyperspectral), and two visual temporal modalities (Optical and SAR). The downstream dataset details are in the supplementary.
@@
Method BigEarthNet (MS)  [50] SoSAT-LCZ42 (MS) [75] TreeSatAI (RGBI) [4] Ship Dataset (SAR Imagery)
@@
Object Detection: We curate instruction-following tasks utilizing three tags i.e., refer, identify, and grounding to perform region-level captioning, referring expressions, and grounded description for object detection datasets from various remote imaging modalities like optical, SAR, and infrared. We include visual grounding [52, 66] datasets for region-level captioning. Following [28], we compute the key attributes in the image such as the object’s category, bounding box, color, relative position, and relative size present in the detection dataset. We present the box as $[x_{min},y_{min},x_{max},y_{max},\theta]$. Here, $(x_{min},y_{min})$ denotes the top left corner point while $(x_{min},y_{min})$ presents the bottom right corner of the bounding box. The angle $\theta$ represents the rotation angle of the bounding box.
@@
EarthDial 73.38 73.34 62.72 62.11 62.11 53.18 72.14 72.14 60.01 63.1 63.1 54.83 68.8 68.8 62.45 61.96 61.96 47.42 64.03 64.03 52.82 63.47 63.47 54.09
Table 6: Comparison of our EarthDial with existing generic and specialized Vision-Language Models (VLMs) for region captioning task across various datasets. R-1, R-L, and MT denotes ROUGE-1, ROUGE-L, and METEOR scores, ZS represents the zero-shot evaluation.
Model HIT UAV [53] (ZS) NWPU VHR 10 [13] (ZS) Swimming Pool Dataset (ZS) UCAS AOD [74] (ZS)
@@
EarthDial 2.61 13.86 28.31 28.06 22.25 17.07 41 27.05 26.35 23.12 1.9 7.4 29.7 29.31 22.77 8.5 34.02 21.17 20.28 13.01
Table 7: Comparison of EarthDial with existing generic and specialized Vision-Language Models (VLMs) on the grounding description task across multiple datasets. R-1, R-L, and MT denotes ROUGE-1, ROUGE-L and METEOR scores. Here ZS means zero-shot evaluation.
@@
Table 9: Performance comparison between our EarthDial and other existing VLMs for visual question answering task (left: RSVQA-LRBEN, right: RSVQA-HRBEN). We report average recall. Comp: Comparison, R/U: Rural/Urban.
@@ 5 Experiments
Here, we discuss the experimental results of the proposed EarthDial across a diverse set of applications, including RGB, multispectral, SAR, infrared, and thermal imagery. Our evaluation covers various tasks such as scene classification, referred object detection, region captioning, grounding descriptions, VQA, image captioning, change detection, methane plume detection and urban heat island.
Scene classification: For zero-shot evaluation, we compare EarthDial with RGB datasets in Tab. 3, whereas BigEarthNet (RGB) was supervised. We also compare multi-spectral (MS) datasets (BigEarthNet, SoSAT-LCZ42), and the RGB-Infrared TreeSatAI dataset as shown in Tab. 4. We notice that EarthDial shows consistent performance gain against existing generic and specialized VLMs. Moreover, EarthDial outperforms temporal scene classification FMoW as well as over xBD test-set 1 for disaster assessment [24] as in Tab. 3.
Object Detection: Following [28], we address three sub-tasks: referred object detection, region captioning, and grounding description. We consider existing generic VLMs like GPT-4o, InternVL2-4B while specialized GeoChat for the comparison. As existing InternVL2 doesn’t provide the rotated bounding boxes, for fair comparison, we finetune the InternVL2 on GeoChat-Instruct and compared it with our EarthDial. Table 5,  6, and  7 depict our EarthDial is a clear winner and consistently outperforms the all other compared VLMs by a large margin. Especially on grounding description tasks, where existing methods struggle to detect/localize the objects, our Earthdial achieves higher mAP. Also, improved results on SAR imagery datasets showcase the multi-modal data processing capability of our EarthDial.
@@
ROUGE-1 ROUGE-L METEOR ROUGE-1 ROUGE-L METEOR ROUGE-1 ROUGE-L METEOR ROUGE-1 ROUGE-L METEOR
@@
EarthDial 31.94 30.66 55.83 33.78 30.47 74.8 30.03 25.69 34.98 18.03 17.42 14.98
Table 10: Comparison of EarthDial with existing generic and specialized Vision-Language Models (VLMs) on the change detection task across multiple datasets.
@@
In the main paper, we report results for Image Classification test Set-1. Here, Tab. 11 presents a summary of the performance results of both generic and specialized VLMs across the sub-tasks discussed above. EarthDial consistently outperforms all other existing VLMs by a significant margin, demonstrating its capability to effectively process temporal data for the desired task.
Additionally, we evaluated our method on QuakeSet [8] for earthquake prediction using SAR imagery. We evaluate the model based on binary classification to determine whether an earthquake event occurred or not between the input SAR imagery. While GPT-4o achieves a classification accuracy of 55.86, our method outperforms it with an accuracy of 57.53.
Multi-modal Data Processing: To showcase the capability of EarthDial in processing multi-modal data, we examine multi-spectral (MS), RGB-infrared, and SAR imagery for classification and referred object detection tasks. Comparative results of the proposed EarthDial and existing GPT-4o are given in Table 4. Our EarthDial outperforms GPT-4o by a significant margin on both multi-spectral, RGBI, and SAR imagery. Significant improvement in the performance highlights the effectiveness of our multi-band fusion strategy.
Urban Heat Island: For UHI, we prompt to classify the underlying region into cooler, mildly hot, and extremely hot regions. We achieve an accuracy of 57.40% in identifying the temperature trends from user-input Landsat8 bands, whereas GPT-4o achieves an accuracy of 22.68% in the same task. This shows the capability of our EarthDial in processing multi-modality data effectively compared to the existing generic GPT-4o.
Methane Plume Classification: We consider STARCOP dataset for the evaluation, which has a four-channel image (RGB + mag1c band). Our EarthDial processes these RGBM bands to predict the presence of methane plumes as ”Yes” or ”No”. We compare our accuracy with the GPT4o model. We simply give RGB and Mag1c bands to the GPT4o and collect the response. With this, GPT4o achieves an accuracy of 40.93%, while our EarDial achieves 77.09% i.e., improves it by 32.16%.
@@ 6 Ablation Study
Pre-training Small Medium Large Single Multiple Data BigEarthNet TreeSat
@@ 7 Conclusion
We present EarthDial, a conversational assistant purpose-built for Earth Observation (EO) data, capable of transforming complex, multi-sensory Earth observations into interactive, and natural language dialogues. EarthDial supports multi-spectral, multi-temporal, and multi-resolution imagery as input and addresses a wide spectrum of remote sensing tasks, including classification, detection, captioning, question answering, visual reasoning, and visual grounding. To enable this versatility, we developed a comprehensive instruction-tuning dataset containing over 11M instruction pairs, encompassing diverse modalities such as RGB, S2, Synthetic Aperture Radar (SAR), Near-Infrared (NIR), and infrared. Additionally, EarthDial excels in handling bi-temporal and multi-temporal sequence analysis, making it highly effective for applications like change detection/disaster assessment. Our experiments across 37 downstream tasks highlight EarthDial’s superior performance over existing generic and domain-specific models, demonstrating its robust generalization capabilities and its potential to set a new standard in EO task automation.
Appendix A Appendix
@@
A.1 EarthDial-Instruct Dataset
@@
Model Swimming Pool Dataset (ZS) Urban Tree Crown Detection [65] (ZS)
GeoChat [28] - 3.1 7.3 1.2 0.6 - 1.8 8.9 2.9 3.1
InternVL2-4B [12] 0.6 6.6 8.9 4.5 0.865 - 3.17 13.41 5.9 3.1
EarthDial-Lora 1.3 2.6 9.45 4.3 0.7 0.2 2.6 9.2 4.1 2.6
EarthDial (Ours) 1.04 7.4 24.90 8.4 1.04 1.1 7.01 25.67 11.13 6.7
@@
Downstream Tasks Image-text Instruction
Though pre-training enhances the generalization capabilities, we also need task-specific fine-tuning with diverse data types to improve downstream performance as shown in Tab. A.2. We curate a large number of instruction-following datasets that include ten diverse downstream tasks: scene classification, object detection, visual question answering, image captioning, change detection, Methane plume detection, tree species classification, local climate zones, urban heat islands, and disaster assessment. It covers seven diverse visual modalities that include Optical, SAR, S2, Infrared, NIR, Landsat8, and Hyperspectral, and two visual temporal modalities (Optical and SAR).
A.2 Ablation on LoRA vs Full Fine-tuning
A.3 Qualitative Analysis
@@
Disaster Assesment QuakeSet [8] train/test SAR User: Do input images present earthquake effects? EarthDial: Yes. User: Could you tell the magnitude the earthquake? EarthDial: 5.58mb.

### v2

@@ Abstract
Automated analysis of vast Earth observation data via interactive Vision-Language Models (VLMs) can unlock new opportunities for environmental monitoring, disaster response, and resource management. Existing generic VLMs do not perform well on Remote Sensing data, while the recent Geo-spatial VLMs remain restricted to a fixed resolution and few sensor modalities. In this paper, we introduce EarthDial, a conversational assistant specifically designed for Earth Observation (EO) data, transforming complex, multi-sensory Earth observations into interactive, natural language dialogues. EarthDial supports multi-spectral, multi-temporal, and multi-resolution imagery, enabling a wide range of remote sensing tasks, including classification, detection, captioning, question answering, visual reasoning, and visual grounding. To achieve this, we introduce an extensive instruction tuning dataset comprising over 11.11M instruction pairs covering RGB, Synthetic Aperture Radar (SAR), and multispectral modalities such as Near-Infrared (NIR) and infrared. Furthermore, EarthDial handles bi-temporal and multi-temporal sequence analysis for applications like change detection. Our extensive experimental results on 44 downstream datasets demonstrate that EarthDial outperforms existing generic and domain-specific models, achieving better generalization across various EO tasks. Our source codes and pre-trained models are at https://github.com/hiyamdebary/EarthDial.
@@ 1 Introduction
  • Experimental results demonstrate that EarthDial performs well in comparison to existing domain-specific VLMs, achieving higher accuracies and better generalization across 44 downstream EO tasks.
@@ 2 Related Work
RS Instruction Datasets: A large number of instruction-following datasets have been introduced to train the RS-VLMs effectively. For example, GeoChat-Instruct [28], SkyEye-968k [67], and RS5M [71] provide extensive RS image-text pairs, supporting instruction-based VLM training on optical data. These existing datasets limit the model’s capabilities across different sensor modalities. On the other hand, MLLMs [70], provide image-text pairs from optical, SAR, and infrared images. Regardless of the large scale, the aforementioned datasets often lack diverse RS applications across various modalities, including multi-resolution, multi-spectral, and multi-temporal RS sensor data. The recently introduced dataset aims to enhance content richness with factual and deceptive questions, enabling VLMs to address a wider range of tasks in the RS domain [46]. However, there is no large-scale unified instruction-following dataset that can encapsulate the distinctive contextual complexities of diverse RS applications across different modalities, multi-resolution, multi-spectral, and multi-temporal RS sensor data. Our work is an effort to bridge this gap with an 11M instruction set to seamlessly integrate multiple earth observation modalities covering diverse spectral and with time-series imaging data for diverse RS applications.
@@ 3.2 Stage-wise Model Training
We utilized the pretrained weights from the previous stage 2 and fine-tuned the MLP and LLM layers, while keeping the ViT parameters frozen from Stage 1. The same ViT+MLP+LLM architecture was employed, with the addition of the data fusion module (Sec. 3.1) to integrate information from multiple spectral channels. This training stage allows the model to handle a broader range of tasks, including land cover classification, species detection, Methane plume detection, UHI and SAR-based ship detection. Training hyperparameter details are given in Table 3.
@@ 4 EarthDial-Instruct Dataset
Pretraining Instruction Data: With EarthDial, our primary aim is to improve generalization performance on diverse downstream tasks, covering a wide range of modalities, multi-resolution, and multi-temporal data. Therefore, we curate high-quality pre-train question-answer (QA) instruction pairs from SkyScript [58] and SatlasPretrain [6] data, which includes Sentinel-2 (S2), Sentinel-1 (SAR), NAIP, and Landsat imagery along with labels. We choose InternLM-XComposer2 [22] for generating instructions using labels. Our data curation process involves filtering to ensure data quality. First, we filter out samples with sparse labels ($<$3). Second, we apply luminance and coverage-based filtering to remove cloudy and low spatial coverage images. Third, we prompt the InternLM-XComposer2 to generate QA instruction pairs based on the key attributes (points, polygons, object category, and position) specified in the inputs and labels. The details can be found in the supplementary material. The curated instruction stats across different imagery sources are presented in Tab. 2.
@@
Stages QA pairs Epoch Train-time GPUs Batchsize Grad. Accu. Lr Decay Adaptive patch sizes
2 1.8M 1 4 (hours) 4A100 80G 2 64 $54e-$ 5 0.05 1 to 6
3 2.4M 1 6 (hours) 4A100 80G 2 64 $54e-$ 5 0.05 1 to 6
Table 3: Details of stages, hyperparams, & training duration
@@
Method BigEarthNet  [50] (MS) SoSAT-LCZ42 [75] (MS) TreeSatAI [4] (RGBI) Ship Dataset (SAR Imagery)
@@
Downstream Tasks Image-text Instruction: While pre-training focuses on enhancing generalization capabilities, we also need task-specific fine-tuning with diverse data types to improve downstream performance. To handle this, we carefully curate a large number of instruction-following datasets that cover ten diverse downstream tasks (e.g., scene classification, object detection, visual question answering, image captioning, change detection, Methane plume detection, tree species classification, local climate zones, urban heat islands, and disaster assessment), six visual modalities (Optical, SAR, S2, Infrared, NIR, and Hyperspectral), and two visual temporal modalities (Optical and SAR). Details are in the supplementary.
@@
Object Detection: We curate instruction-following tasks utilizing three tags i.e., refer, identify, and grounding to perform region-level captioning, referring expressions, and grounded description for object detection datasets from various remote imaging modalities like optical, SAR, and infrared. We include visual grounding [52, 66] datasets for region-level captioning. Following [28], we compute the key attributes in the image such as the object’s category, bounding box, color, relative position, and relative size present in the detection dataset. We present the box as $[x_{min},y_{min},x_{max},y_{max},\theta]$. Here, $(x_{min},y_{min})$ denotes the top left corner point while $(x_{max},y_{max})$ presents the bottom right corner of the bounding box. The angle $\theta$ represents the rotation angle of the bounding box.
@@
EarthDial 73.38 73.34 62.72 61.83 61.83 52.80 72.14 72.14 60.01 63.1 63.1 54.83 68.8 68.8 62.45 61.96 61.96 47.42 64.03 64.03 52.82 63.47 63.47 54.09
Table 7: Comparison of our EarthDial with existing generic and specialized VLMs for region captioning task across various datasets.
Model HIT UAV [53] (ZS) NWPU VHR 10 [13] (ZS) Swimming Pool Dataset (ZS) UCAS AOD [74] (Zero-Shot)
@@
EarthDial 2.61 13.86 28.31 28.06 22.25 17.07 41.00 27.05 26.35 23.12 1.9 7.4 29.7 29.31 22.77 8.5 34.02 21.17 20.28 13.01
Table 8: Comparison of EarthDial with existing generic and specialized VLMs on the grounding description task across multiple datasets.
@@
Table 10: Comparison of EarthDial with existing VLMs for visual question answering task (left: RSVQA-LRBEN, right: RSVQA-HRBEN). Comp: Comparison, R/U: Rural/Urban.
@@ 5 Experiments
Here, we discuss the experimental results of the proposed EarthDial across a diverse set of applications, including RGB, multispectral, SAR, infrared, and thermal imagery. Our evaluation covers various tasks such as scene classification, referred object detection, region captioning, grounding descriptions, VQA, image captioning, change detection, and methane plume detection.
Scene classification: For zero-shot evaluation, we compare EarthDial with RGB datasets in Tab. 4, whereas BigEarthNet (RGB), xBD Set 1, and fMoW are supervised. We also compare multi-spectral (MS) datasets (BigEarthNet, SoSAT-LCZ42), and the RGB-Infrared TreeSatAI dataset as shown in Tab. 5. We notice that EarthDial shows consistent performance gain against existing generic and specialized VLMs. Moreover, EarthDial outperforms temporal scene classification FMoW as well as over xBD test-set 1 for disaster assessment [24] as in Tab. 4. Further results are in a supplementary document.
Object Detection: Following [28], we address three sub-tasks: referred object detection, region captioning, and grounding description. We consider existing generic VLMs like GPT-4o, InternVL2-4B while specialized GeoChat for the comparison. As existing InternVL2 doesn’t provide the rotated bounding boxes, for fair comparison, we finetune the InternVL2 on GeoChat-Instruct and compared it with our EarthDial. Table 6,  7, and  8 depict our EarthDial is a clear winner and consistently outperforms the all other compared VLMs by a large margin. Especially on grounding description tasks, where existing methods struggle to detect/localize the objects, our Earthdial achieves higher mAP. Also, improved results on SAR imagery datasets showcase the multi-modal data processing capability of our EarthDial.
@@
R-1 R-L MT R-1 R-L MT R-1 R-L MT R-1 R-L MT
@@
EarthDial 31.94 30.66 55.83 33.78 30.47 74.8 28.16 24.03 33.56 18.03 17.42 14.98
Table 11: Comparison of EarthDial with existing generic and specialized VLMs on the change detection task.
@@
In the main paper, we report results for Image Classification test Set-1. Here, Tab. 12 presents a summary of the performance results of both generic and specialized VLMs across the sub-tasks discussed above. EarthDial consistently outperforms all other existing VLMs by a significant margin, demonstrating its capability to effectively process temporal data for the desired task. Tab. 12 presents a summary of the performance results of both generic and specialized VLMs across the sub-tasks discussed above. EarthDial consistently outperforms all other existing VLMs by a significant margin, demonstrating its capability to effectively process temporal data for the desired task.R-1, R-L, MT denotes ROUGE-1, ROUGE-L, METEOR scores Additionally, we evaluated our method on QuakeSet [8] for earthquake prediction using SAR imagery. We evaluate the model based on binary classification to determine whether an earthquake event occurred or not between the input SAR imagery. While GPT-4o achieves a classification accuracy of 55.86, our method outperforms it with an accuracy of 57.53.
Multi-modal Data Processing: To showcase the capability of EarthDial in processing multi-modal data, we examine multi-spectral (MS), RGB-infrared, and SAR imagery for classification and referred object detection tasks. Comparative results of the proposed EarthDial and existing GPT-4o are given in Tab. 5. Our EarthDial outperforms GPT-4o by a significant margin on both multi-spectral, RGBI, and SAR imagery. Significant improvement in the performance highlights the effectiveness of our multi-band fusion strategy.
Urban Heat Island: For UHI, we prompt to classify the underlying region into cooler, mildly hot, and extremely hot regions. We achieve an accuracy of 56.77% in identifying the temperature trends from user-input Landsat8 bands, whereas GPT-4o achieves an accuracy of 22.68% in the same task. This shows the capability of our EarthDial in processing multi-modality data effectively compared to the existing generic GPT-4o.
Methane Plume Classification: We consider STARCOP dataset for the evaluation, which has a four-channel image (RGB + mag1c band). Our EarthDial processes these RGBM bands to predict the presence of methane plumes as "Yes" or "No". We compare our accuracy with the GPT4o model. We simply give RGB and Mag1c bands to the GPT4o and collect the response. With this, GPT4o achieves an accuracy of 40.93%, while our EarDial achieves 77.09% i.e., improves it by 32.16%.
@@ 6 Ablation Study
AHR Pre-training Small Medium Large Single Multiple Data BigEarthNet TreeSat
@@ 7 Conclusion
We present EarthDial, a conversational assistant purpose-built for Earth Observation (EO) data, capable of transforming complex, multi-sensory Earth observations into interactive, and natural language dialogues. EarthDial supports multi-spectral, multi-temporal, and multi-resolution imagery as input and addresses a wide spectrum of remote sensing tasks, including classification, detection, captioning, question answering, visual reasoning, and visual grounding. To enable this versatility, we developed a comprehensive instruction-tuning dataset containing over 11M instruction pairs, encompassing diverse modalities such as RGB, S2, Synthetic Aperture Radar (SAR), Near-Infrared (NIR), and infrared. Additionally, EarthDial excels in handling bi-temporal and multi-temporal sequence analysis, making it highly effective for applications like change detection/disaster assessment. Our experiments across 44 downstream tasks highlight EarthDial’s superior performance over existing generic and domain-specific models, demonstrating its robust generalization capabilities and its potential to set a new standard in EO task automation.
Appendix
@@
Appendix A EarthDial-Instruct Dataset
@@
Disaster Assesment QuakeSet [8] train/val/test SAR User: Do input images present earthquake effects? EarthDial: Yes. User: Could you tell the magnitude the earthquake? EarthDial: 5.58mb.
@@
Downstream Tasks Image-text Instruction
Though pre-training enhances the generalization capabilities, we also need task-specific fine-tuning with diverse data types to improve downstream performance as shown in Tab. A.1 and Fig. A.2. We curate a large number of instruction-following datasets that include ten diverse downstream tasks: scene classification, object detection, visual question answering, image captioning, change detection, Methane plume detection, tree species classification, local climate zones, urban heat islands, and disaster assessment. It covers seven diverse visual modalities that include Optical, SAR, S2, Infrared, NIR, Landsat8, and Hyperspectral, and two visual temporal modalities (Optical and SAR).
Model Swimming Pool Dataset (ZS) Urban Tree Crown Detection [65] (ZS)
GeoChat [28] - 3.1 7.3 1.2 0.6 - 1.8 8.9 2.9 3.1
InternVL2-4B [12] 0.6 6.6 8.9 4.5 0.865 - 3.17 13.41 5.9 3.1
EarthDial-Lora 1.3 2.6 9.45 4.3 0.7 0.2 2.6 9.2 4.1 2.6
EarthDial (Ours) 1.04 7.4 24.90 8.4 1.04 1.1 7.01 25.67 11.13 6.7
Figure A.2: Illustration of our versatile EarthDial model that performs across multi-modalities, multi-resolution, multispectral, and multitemporal data from diverse remote sensing applications. EarthDial extends its capabilities to a range of tasks such as scene classification, image/region-captioning, referring expression, VQA, referring expression, object detection, temporal change/disaster detection, Methane plume detection, tree species classification, UHI, and LCZs detection across multi-modalities, multi-resolution remote sensing data.
Appendix B Ablation on LoRA vs Full Fine-tuning
B.1 Qualitative Analysis: